OpenStudy (anonymous):

Solve by elimination. 2x - 5y = 11 4x + 10y = 18 HELP

OpenStudy (ash2326):

@algebra2sucks did you try this? do you know to use elimination method?

OpenStudy (anonymous):

not 2 sure of it i'm not that good in algebra 2

OpenStudy (ash2326):

Could you tell how did you approach?

OpenStudy (anonymous):

what ui do it as 2 combing the questions or solving each one for its own?

OpenStudy (ash2326):

In elimination method. To solve a pair of linear equations, we'd eliminate one of the two variable. First step: Multiply both or one of the equations by some appropriate constants Second step: Add or subtract the equations to eliminate on of the variables. I'll illustrate this methodology here

OpenStudy (ash2326):

Are you here?

OpenStudy (anonymous):

[2x - 5y = 11 ] * 2 ==> 4x - 10y = 22 4x + 10y = 18 =================== 4x - 10y = 22 4x + 10y = 18 -------------add both equations

OpenStudy (anonymous):

so do i add 22 and 18? or 4x and 10y or both and make it into one equation?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

40 + 14?

OpenStudy (anonymous):

@surdawi or @ash2326 is that right?

OpenStudy (ash2326):

\[4x - 10y = 22 \] \[4x + 10y = 18 \] Add all the terms, remember that x can only be combined with an x not a y or 22 or 18 \[(4x+4x)+-10y+10y=22+18\] what would you get?

OpenStudy (anonymous):

4x + 4x = 8x -10y + 10y = 0y 22 + 18 = 40

OpenStudy (ash2326):

\[8x+0y=40\]

OpenStudy (anonymous):

so do i have to make it more simple to get one answer?

OpenStudy (anonymous):

8x = 40 now?

OpenStudy (anonymous):

would i divide the two?

OpenStudy (anonymous):

yes, what do u get

OpenStudy (anonymous):

5x?

OpenStudy (anonymous):

show me how?

OpenStudy (anonymous):

i'm not sure i did it on a calculater when i did it both ways first i got 0.2

OpenStudy (anonymous):

what numbers did u plug in

OpenStudy (anonymous):

i put 40/8 and got 5 when i put 8/40 i got 0.2

OpenStudy (anonymous):

8x = 40 /8 /8 x = 5

OpenStudy (anonymous):

thank you for all ur help